OpenStudy (anonymous):

(y+3)^(2/3)-2(y+3)^(1/3)-3=0

OpenStudy (anonymous):

Let (y+3)^1/3 = a a^2 -2a -3 =0 (a-3)(a+1) =0 a =3, a = -1 (y+3)^1/3 = 3 or (y+3)^1/3 = -1 Take cubes on both side y+3 = 27 or y +3 = -1 y = 24 or y = -4

OpenStudy (anonymous):

y =24